soybean is 16% protein; cornmeal is 8% protein. How many pounds of each should be mixed together in order to get 320 lb mixture that is 11% protein. Can you help me on this????
Answer by mananth(16946)

percent ---------------- quantity
Soyabean 16 ---------------- x
Corn 8 ---------------- 320-x
Total 11.00 ---------------- 320
...
16x+8(320-x)=11* 320
16x+2560 -8x=3520
16x-8x =3520-2560
8x=960
/8
x=120 pounds Soyabean
200 pounds of corn meal